# NGO Activism:

### v7

@@ 2 Data
We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year). Appendix Table A1 reports descriptive statistics. Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once. Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign. Panel (b) documents substantial variation across targeted firms in size and news coverage. Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations. 53% of the campaigns are on social causes, with the remaining being environmental. We return to the remaining panels of Table A1 in Section 4, after first providing background on AGM dates.
3 Institutional Background on AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B1 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^2] Results remain robust excluding the covid period (2020–2021).
While external events may shape NGOs’ incentives to campaign, they do not alter the timing of AGMs. This stability makes AGMs a useful benchmark: NGOs can choose to react immediately to a negative event concerning a firm, or alternatively wait and align their campaign with its AGM, when the firm is most exposed to scrutiny, as shown in Panel (b) of Figure 1, which shows that news coverage spikes in AGM weeks. Consistent with this view, we show in the Appendix that AGM-timed campaigns generate more negative news coverage than those launched at other times (Appendix Figure B2). Taken together, these patterns highlight why AGMs provide a natural focal point to study the strategic trade-off between maximizing visibility and exerting direct influence.
4 Timing Campaigns for Visibility
Having shown that NGO campaigns cluster around AGM dates and that they increase a target’s negative media coverage, we next ask whether such timing enhances NGOs’ visibility. We use Google search activity as a proxy for NGO visibility and estimate event-study regressions comparing campaigns launched on the AGM date with those launched at other times:
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{c}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$ (1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\mathbf{1}_{\{M_{c,m,y}=\tau\}}\right)$
$\quad+\alpha_{c,y}+\alpha_{m,y}+u_{g,m,y}.$
Because NGOs typically campaign multiple times, we follow Baker et al. (2022a) and stack all campaign events. Each event is denoted $c(i,n,g,y,m)$, where campaign $c(\cdot)$ is identified by cause $i$ (e.g., pollution), NGO $g$ (e.g., Greenpeace), target firm $n$ (e.g., Shell), calendar year $y$, and month $m$. We then analyze variation in the Google Trends score for NGO $g$ within a five‐month window around the start of $g$’s campaign; this window may spill into years y1y-1 or $y+1$ if the campaign occurs early or late in year $y$.
The main explanatory variables are event-time dummies, $\mathbf{1}_{\{M_{c,m,y}=\tau\}}$, equal to one if month $m$ is exactly $\tau$ months from campaign $c$, with $\tau\in[-5,5]$. The omitted category is the month immediately preceding the campaign ($\tau=-1$). These indicators are interacted with $\text{AGM Campaign}_{c}$, a dummy equal to one if campaign $c$ was launched on the target firm’s AGM date. Thus $\beta^{AGM}_{\tau}$ estimates the increase in visibility for NGO campaigning on an AGM date, over the direct effect of campaigning at any date, $\beta^{Non-AGM}$.
The remaining terms include fixed effects at the NGO–firm–year level ($\alpha_{c,y}$), which absorb all variation across NGO-firm pairs over years, as well as month-by-year fixed effects ($\alpha_{m,y}$). Our main coefficients of interest ($\beta^{AGM}_{\tau}$) are therefore identified from within-pair variation across months. Standard errors are clustered at the NGO level.
Under the hypothesis that AGM campaigns affect NGO visibility differently depending on their reputation, we estimate (1) separately for an NGO’s first few campaigns, when its recognition is still limited, and for later ones. Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and subsequent campaigns (red triangles). The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months. For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established. These results are estimated within NGOs: early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
Figure 2: Increased NGO visibility around an AGM campaign
(a) The first AGM campaigns bring visibility
(b) No evidence of Selection
Note: Panel (a) plots the estimated $\beta^{AGM}$ coefficients from (1). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level. Panel (b) reports the estimated coefficients of the regression of an indicator that an NGO campaign was launched on the target firm’s AGM date on the displayed (standardized) variables, and year fixed effects. Error bars shows 90% (shaded) and 95% confidence intervals using standard errors that are clustered by firm.
We show that these results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with substantially larger gains for early campaigns.
Third, although we find no evidence of pre-trends, NGOs might still select targets strategically to raise visibility—for example, by focusing on firms that are more vulnerable to pressure. Panel (b) of Figure 2 addresses this concern by regressing an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the cause of the campaign, with year fixed effects and firm-clustered standard errors. [^3] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), the firm’s Environmental and Social scores (Env. score, Soc. score), how negative the campaign is towards the firm (Campaign sentiment), and an indicator for the campaign cause being related to an environmental topic (Env. campaign). Note that no coefficient is significantly different from zero. We find that a one standard deviation increase in any of these variables moves the likelihood of an AGM campaign by less than 5% of a standard deviation, indicating little selection on observables.
Further evidence on potential selection based on the number of past NGO campaigns is presented in Table A1. Panel (c) compares firms targeted by NGOs that had previously campaigned on a target’s AGM date with those that had not, finding no significant differences across the same variables analyzed in Panel (b) of Figure 2. This suggests that NGOs do not appear to select targets based on their past campaigning strategies. Panel (d) corroborates this result by regressing firm characteristics on the number of past NGO campaigns and showing no systematic relationship once fixed effects are included. The only consistent pattern is that NGOs increasingly focus on environmental rather than social issues over time, reflecting their growing salience.
Thus, the evidence presented so far points to:
Result 1.
Visibility gains from AGM‐day campaigns are greatest during an NGO’s initial campaigns.
For clarity, each of the next sections highlights its main findings in a similar manner.
4.1 Stakeholders’ Reactions
We next examine how campaign timing affects stakeholders and firms’ ESG outcomes.
4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022b). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix D describes the matching procedure.
For each campaign in our sample, we examine the probability that the next AGM includes a related shareholder proposal, depending on the campaign’s timing. To this end, we recalendarize time so that month 0 corresponds to the AGM month for each firm. As discussed in Section 3, AGM months are stable over time; thus, this recentering does not alter the distribution of months across years or risk double counting. [^4] All subsequent analyses use the same recalendarization. Months –5 and +5 are defined as 45-day intervals to ensure the recalendarized year covers the entire calendar. We estimate the following regression:
$+\alpha_{i,s(n),g,m,y}+\gamma_{n}+u_{i,n,y}$
For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause. Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$. Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$ because of SEC rules.
Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across firms, industries, causes, and NGOs over time. We cluster standard errors by firm.
Figure B3 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. This timing suggests that campaigns around an AGM may spur proposals at the next one, possibly due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partners to immediately gain broad support on the proposal, without having to wait a year to advocate for change at the next AGM.
4.1.2 Consumers
Drawing from quarterly data from Compustat, we investigate the responses of consumers to the timing of NGO campaigns by studying the relation between the target’s “Sales to Asset ratio” and the distance between the campaign and the target’s AGM through
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n,y}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n,y}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm-by-year level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns. [^5] Unreported results show also no difference between early and late campaigns.
4.2 Does AGM-Targeting Create Positive Change?
If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2). However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year. [^6] Acknowledging the limits of ESG scores as measures of social and environmental impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), we use MSCI to mitigate signal risk: its ratings are more time-consistent and widely used by sustainable investors (Berg et al., 2021, 2024). By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
This finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
Result 2.
NGO campaigns on AGM dates spur shareholders to place related proposals at future AGMs of the targeted firm, but they produce no contemporaneous improvements in the firm’s E&S scores.
5 Timing Campaigns to Influence AGM Outcomes
Early campaigns could benefit an NGO aiming to influence current AGM outcomes. Since most votes are cast by proxy before the AGM, targeting the AGM date may limit an NGO’s ability to sway a significant share of the votes. For instance, over 85% of vote-eligible shares for Microsoft’s 2023 AGM were cast by proxy before the meeting, indicating that campaigning around AGM dates may have little impact on shareholders’ votes (Microsoft, 2023).
We define a proposal as successful if either it is withdrawn by the shareholder before the AGM or it secures at least $50\%+1$ vote at the AGM. We first look at the influence of NGO campaigns on the success of related shareholder proposals using the following specification on the whole sample of shareholder proposals:
$\text{Success}_{i,n,m,y}=\sumop\displaylimits^{0}_{\begin{subarray}{c}\tau=-6\end{subarray}}\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+\tau,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $\tau$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year yy—i.e., $m\in[-11,-7]$, as campaigns launched during that period are more likely to be in reaction to the previous year’s AGM results. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
(a) Proposal accepted or withdrawn
(b) Vote share in favor of proposal
To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (4) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable. [^7] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and proxy advisors such as ISS (“For,” “Against,” or “Abstain”). Panel (a) of Figure 3 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Panel (a), Panel (b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011). This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
The evidence in this section points to:
Result 3.
Pre-AGM campaigns raise shareholder support and proposal success, while AGM-day campaigns have little effect.
6 Exposure vs. Influence
Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence. In this section, we examine how NGOs navigate this tradeoff as their visibility increases. We first study whether greater past exposure leads NGOs to shift campaign timing from the AGM date to earlier months. We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
6.1 The Timing of Past and Current Campaigns
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
$\quad+\alpha_{t(\cdot),y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$. $\alpha_{t(\cdot),y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively. Identification thus comes from NGOs launching multiple campaigns in the same year. [^8] Appendix Figure B6 shows similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
6.2 Visibility and Campaign Timing
We next use Google Trends scores to proxy for an NGO’s reputation. We compute the average search score for NGO $g$ in year $y$ relative to Greenpeace in year y1y-1, discretized into deciles and regress:
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score}_{g,y-1}=\tau\!\right\}}\right)$ (6)
$\quad+\alpha_{t(\cdot)}+\alpha_{y}+u_{c(i,n,g,y,m)},$
which is similar to model (6) with $\text{Relative Google Score}_{g,y-1}$ as a measure of NGO visibility. $\alpha_{t(\cdot)}$ and $\alpha_{y}$ are topic and year fixed effects. Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
Result 4.
NGOs initially target AGM dates to gain visibility. As their exposure increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
Figure 4: NGO reputation and campaign timing
(a) Later campaigns pre-date the target’s AGM date
(b) More visible NGOs campaign earlier
6.3 NGO Reputation and AGM Influence
We now ask whether NGO reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
Table 1: NGO’s strategy and support for related shareholder proposals
Dependent Variable: Proposal Passes (0/1)
Model: (1) (2) (3) (4)
Campaign on Same Topic -0.0003 -0.013 -0.005 -0.001
(0.026) (0.028) (0.025) (0.030)
# Past Campaigns -0.034
Campaign on Same Topic $\times$ # Past Campaigns 0.078∗∗∗
# Past Campaigns squared 0.008∗∗
Campaign on Same Topic $\times$ # Past Campaigns squared -0.015∗∗
# Past Non-AGM Campaigns 0.0008
Campaign on Same Topic $\times$ # Past Non-AGM Campaigns -0.011
# Past AGM Campaigns 0.005 -0.034
(0.020) (0.023)
Campaign on Same Topic $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
(0.026) (0.032)
# Past AGM Campaigns squared 0.010∗∗
Campaign on Same Topic $\times$ # Past AGM Campaigns squared -0.013∗∗
Campaign on Same Topic $\times$ Google Search Popularity 0.071
Campaign on Same Topic $\times$ Google Search Popularity square -0.021
Fixed-effects
CUSIP x Year
NGO x Year
Topic x Month-Year
Observations 36,543 35,778 36,543 33,142
R2 0.61894 0.61698 0.61995 0.62078
*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
  • Note: This table reports the estimated coefficients from (7) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Topic. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Topic. In Column 4, we replace # Past Campaigns from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has already campaigned ($\beta_{3}>0$). A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average. The regression also includes the squared term of “# $\text{\# Past Campaigns}_{g,y-1}$,” showing decreasing marginal returns of campaigning.
We unpack the mechanism behind these coefficient estimates by distinguishing between prior AGM campaigns and non-AGM campaigns in Column (2). We find that only the former contribute to proposal success. In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
Column (3) tests for diminishing returns by including squared terms. The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate. Importantly, the combined effect remains non-negative even for the most active NGOs. [^9] The marginal effect of AGM campaigning never turns negative in our sample since no firm launches more than five AGM-date campaigns, so our proxy for reputation does not imply adverse returns.
We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends. Specifically, Column (4) replaces “# $\text{\# Past Campaigns}_{g,y-1}$” by “Google Search Popularityg,y-1,” which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.101010The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects. The interaction between this measure and campaigning on the same cause is positive, with a concave shape. Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
These results support the following result:
Result 5.
Targeting AGMs builds NGO reputation and increases their ability to raise the support of related shareholder proposals, but with diminishing returns.
7 Discussion and Conclusion
Examining NGO campaign timing, we show that not just the decision to intervene but also its timing can materially affect firm behavior and shape the course of future activism. Over time, NGOs shift from visibility-driven exposure tactics to those designed to influence corporate strategies, highlighting how the temporal dimension of their actions channels pressure that leads firms to internalize social and environmental externalities.
We first show that NGOs’ early campaigns often coincide with the target’s AGM date, when media attention is highest, and that such timing significantly increases their visibility (Result 1). These visibility gains are concentrated in NGOs’ first few campaigns, indicating that AGM-day actions serve primarily to build name recognition rather than to influence firms directly. By contrast, later AGM-day campaigns have no immediate effect on firms’ E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
Hence, campaigning before the AGM offers direct advantages. When an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of success—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3). Given the typically low baseline support for such proposals, early NGO interventions can effectively discipline managers who must contend with heightened shareholder pressure (Levit and Malenko, 2011). Importantly, all analyses include NGO-by-year and cause-by-year fixed effects, thereby accounting for time-varying confounders at both the NGO level—such as experience or reputation—and the cause level—such as specific news events. This ensures that globally renowned NGOs (e.g., PETA) do not drive the results. Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
Our results suggest that NGOs face a trade-off that they navigate by adapting their strategies over time. NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4).
As a result, having targeted AGMs in the past increases the likelihood that an NGO can help a shareholder proposal pass related to its campaign. However, repeatedly targeting AGMs leads to weaker influence over firm behavior (Result 5).
One interpretation is that NGOs, having already demonstrated commitment to their cause, face weaker incentives to keep signaling through costly AGM-day campaigns. In this view, declining returns simply reflect a reduced need for visibility once reputational capital is established.
Yet the evidence points to a different mechanism. Within the same NGO, AGM-day campaigns progressively lose their marginal effect on media attention (Figure 2) and their ability to damage a firm’s reputation (Appendix Figure B10), suggesting that the audience’s response—rather than the NGO’s effort—changes over time. As NGOs become recognized and credible actors, stakeholders interpret their campaigns differently: what initially signals toughness later signals redundancy or routine. Established NGOs are expected to engage in more constructive dialogue, such as supporting shareholder proposals, rather than resorting to public confrontation. Consequently, the same action produces less reputational harm for firms and fewer visibility gains for NGOs.
The diminishing returns to AGM targeting thus stem from evolving stakeholder expectations and a changing perception of NGO legitimacy, rather than from reduced internal pressure to seek exposure. [^11] This interpretation does not rule out signaling motives altogether. Early AGM-day campaigns are consistent with “reputation-for-toughness” models, in which actors bear short-term costs to establish credibility that can be leveraged in future interactions (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex) or experience in managing AGM-related information is accounted for through NGO-by-year fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs without an established reputation coordinate stakeholders over time by campaigning on AGM dates, drawing attention to issues that spur future engagement. Once reputable, they coordinate stakeholders within the same AGM cycle by supporting existing shareholder proposals and directly shaping firm decisions. Thus, NGOs act as a coordination device throughout their lifecycle, with the form of coordination evolving with their perceived toughness and credibility. In this way, NGOs’ incentives shape their strategies and, ultimately, the social and environmental impact of firms.
We also acknowledge that NGOs increasingly engage not only in campaigns but also in collaborations with corporations (Dahan et al., 2010). Future research could examine when confrontation shifts to cooperation and how such collaboration reshapes the balance between visibility, credibility, and influence.
[…]
Appendix A Omitted Tables
Table A1: Summary Statistics
@@ Appendix A Omitted Tables
Panel c: Selection—Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past

### v8

@@ 3 Data
We link Sigwatch, ISS, and RavenPack data using standard identifiers (CUSIP, ISIN, year). Table 1 reports descriptive statistics. Panel (a) shows that, on average, NGOs launch 3.2 campaigns during the sample, 24% of which are on the target AGM date, though over half of NGOs appear only once. Most campaigns focus on social rather than environmental issues, but the distribution is highly skewed: a small set of NGOs accounts for a disproportionate share of activity, while the median NGO runs only a single campaign. Panel (b) documents substantial variation across targeted firms in size and news coverage. Targets range from very large firms with extensive media presence to smaller firms with limited coverage, suggesting that NGOs are not exclusively focused on high-visibility corporations. 53% of the campaigns are on social causes, with the remaining being environmental. We return to the remaining panels of Table 1 in Section 5, after first providing background on AGM dates.
Table 1: Summary Statistics
@@
Form 990-filing U.S. NGO 0.15 0 0 0 1
     – Number of campaigns 10.19 1 3 14 33
     – Campaigns on Env. topic 4.36 0 1 4 19
     – Campaigns on Soc. topic 2.83 0 1 4 12
     – On AGM date 0.28 0 0 0 1
Campaign sentiment -1.75 -2 -2 -2 -1
NGO Total Assets (USD mn) 50.02 2.80 7.95 53.91 248.02
Yearly donations received (USD mn) 38.03 3.18 10.13 59.14 143.09
NGO age (years) 42.79 21 32 47 126
@@
4 Empirical Strategy
Our empirical strategy exploits the timing of NGO campaigns relative to AGMs, leveraging the fact that AGM dates are largely exogenous and stable within firms over time due to SEC regulations. Rule 14a-8 requires shareholder proposals to be filed at least 120 days before the proxy statement, effectively anchoring AGM dates across years (glac2014influence; fioretti2024shared). Figure 2 illustrates this stability by comparing each firm’s AGM month across consecutive years. Most observations lie along the diagonal, indicating strong persistence of the AGM month.
This predictability is central to our identification strategy. Because AGM dates are fixed well in advance and rarely change, NGOs face an exogenous schedule of heightened public attention: they can either react immediately to external events or strategically align their campaigns with these predetermined meetings, when firms are most exposed to shareholder and media scrutiny. In this section, we exploit this feature to study how campaigning around AGMs affects NGOs’ visibility (Section 5), stakeholder reactions (Section 5.1), and firms’ subsequent outcomes (Section 6.1).
Figure 2: Persistence of AGM month over time
Note: The figure plots the percentage of AGMs held in each month (y-axis) given the month of the previous year’s AGM (x-axis). Larger and darker circles indicate a higher proportion of meetings. Data cover 38,471 meetings of 5,745 U.S. companies from 2010–2020.
4.1 Target Selection
A potential concern is that NGOs might select target firms strategically to maximize either visibility or credibility rather than choosing campaign timing exogenously. For instance, NGOs could preferentially target firms that are already more visible, frequently covered in the media, or more responsive to shareholder pressure. Table 2 addresses this concern by testing for selection on observables o
Panels (a) and (b) examine whether NGOs with prior AGM campaigns systematically target different types of firms. Panel (a) compares average firm characteristics across NGOs with and without previous AGM campaigns. In addition, Appendix Figure B1 regresses an indicator for AGM-timed campaigns on firms’ financial and ESG characteristics and the campaign’s cause, controlling for year fixed effects and clustering standard errors at the firm level. [^6] The independent variables include the target firm’s market capitalization (Market Cap), yearly sales (Sales), total assets (Total Assets), number of news over the past 12 months (News count), environmental and social scores (Env. score, Soc. score), campaign sentiment, and an indicator for whether the campaign concerns environmental topics (Env. campaign). None of these variables significantly predicts AGM targeting. A one–standard-deviation increase in any firm characteristic shifts the likelihood of an AGM campaign by less than 5 percent of a standard deviation, indicating little selection on observables regarding on whether target firms of NGOs targeting AGM campaigns systematically differ from other NGOs.
Panel (b) further regresses target characteristics on the number of an NGO’s past campaigns (the results are similar for past AGM campaigns). We include standard balance-sheet variables (market capitalization, total assets, and sales) as well as media and ESG indicators (news coverage, environmental and social scores, and topic dummies). None of these coefficients is economically and statistically significant, except for a mild increase in environmental campaigns over time, consistent with a post–Paris Agreement shift in NGO focus toward environmental issues. Panels (c) and (d) turn to NGO-level selection. Panel (c) focuses on 541 U.S. NGOs with balance-sheet data and regresses firm characteristics on NGO size, proxied by deciles of total assets, while Panel (d) correlates NGO financial ratios – assets, donations-to-assets, and expenses-to-assets – with the number of past campaigns. In both analyses, coefficients are small and statistically insignificant, showing no systematic link between NGO scale or financial capacity and target-firm characteristics.
Taken together, these results indicate that neither NGO nor firm characteristics predict AGM-timed activism. The evidence supports our interpretation that campaign timing reflects strategic decisions by NGOs rather than endogenous target selection. However, while our analysis rules out selection on observables, unobserved factors may still affect campaign timing – a possibility we address in Section 8, where we exploit quasi-exogenous variation in campaign dates.
Table 2: Selection on Observables
Panel a: Avg. target characteristics’ based on whether the campaigning NGO campaigned on AGM dates in the past